# Unmasking Trees for Tabular Data

### v2

@@ 3.3 Evaluating BaltoBot on synthetic probabilistic prediction case studies
We compare our approach with Treeffuser (Beltran-Velez et al., 2024) on the “wave” synthetic dataset from Treeffuser (Beltran-Velez et al., 2024), which as shown in Figure 3 is nonlinear, multimodal, heteroskedastic, and heavy-tailed. On the raw probabilistic predictions in Figure 3(A), we see that BaltoBot is (by visual inspection) able to model the conditional distribution as well as Treeffuser; BaltoBoTabPFN performs slightly worse. Yet this case study illustrates the two advantages of BaltoBot. First, in Figure 3(B) we show the runtime of the different methods: training, sampling, and total. To train on 5000 samples, Treeffuser took 1.1s and BaltoBot took 2.6s; but to generate 5000 samples, Treeffuser took 5.0s while BaltoBot took 0.72s. Second, BaltoBot offers the ability to estimate a closed-form probability density function (pdf) of the predictive distribution in Figure 3(C); in contrast, Treeffuser can only sample from the predictive distribution.
@@
4 Discussion
@@
Diffusion modeling has recently gained popularity in tabular ML (Zheng and Charoenphakdee, 2022; Jolicoeur-Martineau et al., 2024b; Beltran-Velez et al., 2024; Kotelnikov et al., 2023). Our proposed approach is an instance of the autoregressive discrete diffusion framework (Hoogeboom et al., 2021), instances of which have shown success in a variety of tasks (Yang, 2019; Austin et al., 2021; Kitouni et al., 2024; Jolicoeur-Martineau et al., 2024a). Yet our results call into question whether diffusion is beneficial for tabular conditional generation, or whether autoregression is sufficient for our setting. It has been observed that diffusion is autoregression in frequency space, progressing from low frequencies to high frequencies, which makes it a good match for image data with its power law spectra (Rissanen et al., 2022; Dieleman, 2024; Stews, 2024). In tabular datasets without this phenomena, we would expect diffusion modeling to be less advantageous. Our success also makes sense in light of the observation that unmasking tends to outperform denoising in self-supervised pretraining (Balestriero and LeCun, 2024).
@@
Why is ForestDiffusion better at vanilla generative modeling, while UnmaskingTrees is better on generation given partially-missing data and especially on imputation? We offer two speculative explanations. First, imputation is a conditional modeling scenario, except that you do not know the partition of the features into input features and output features a priori. One could address imputation by learning all possible $2^{D}$ conditional distributions, but this is impractical for large $D$, so one would prefer to learn a single joint distribution. Both autoregression and diffusion are ways of learning a joint distribution; because autoregression does so by learning conditional distributions, it is more suited to the conditional modeling imputation setting. Second, for missing data, diffusion has a train-inference gap: during training, observed features begin the reverse process from $\mathcal{N}(0,1)$; during inference for imputation, observed features begin the reverse process at their actual values. On the other hand, the advantages of diffusion modeling (no quantization error, holistic generation) give it superiority when these problems can be avoided.
@@
We also note that MissForest (Stekhoven and Bühlmann, 2012), hailing from statistical literature on multiple imputation, has yet to be fully dethroned. Future progress in tabular conditional generation may require going back to the well of this traditional literature. As one example, we observe that MissForest exploits feature missingness fraction information, but we are not aware of any “machine learning” approaches which do so. The statistical literature has also previously explored the value of conditional modeling for joint modeling (Gelman and Raghunathan, 2001; Liu et al., 2014; Kropko et al., 2014). Indeed, our UnmaskingTrees approach, and all autoregressive modeling, is presaged by the full-mechanism bootstrap (Efron, 1994).
Finally, we observe where randomness enters into our generation process. Flow-matching injects randomness solely at the beginning of the reverse process via Gaussian sampling, whereas diffusion models inject randomness both at the beginning and during the reverse process. In contrast, because our method starts with a fully-masked sample, it injects randomness gradually during the generation process. First, we randomly generate the order over features for unmasking. Second, we do not “greedily decode” to the most likely leaf in the meta-tree, but instead sample according to predicted probabilities. Third, for continuous features, having sampled a particular meta-tree leaf bin, we sample from within the bin, treating it as a uniform distribution.
@@ 5 Conclusions
We show that tree-based autoregressive unmasking is a strong, simple baseline for tabular data. We recommend UnmaskingTrees for tabular imputation, especially when downstream predictions are the goal, and for generation on datasets with missingness. We also offer BaltoBot as a fast, flexible method for computing predictive distributions.
Appendix A Ablation experiments with imputation
We additionally run UnmaskingTrees without BaltoBot, and instead with vanilla quantization using k-Means [Lloyd, 1982] and KDI [McCarter, 2023]. Average ranks (shown in Table 4) and raw scores (shown in Table 5) demonstrate that UnmaskingTrees on its own improves upon Forest-VP’s diffusion approach. We also see that KDI quantization contributes to improvement beyond k-Means, and that BaltoBot yields even further improvement.
KNN 6.8 (0.6) 7.8 (0.6) 6 (0.4) 6.1 (0.5) 10.4 (0) 8.2 (1.3) 7 (1.5) 7.5 (1.5) 6.5 (0.8)
ICE 8.3 (0.5) 5.8 (0.5) 8.5 (0.6) 8.8 (0.5) 1.9 (0.4) 8 (1.1) 9 (0.6) 7.2 (1.1) 6.4 (0.8)
MissForest 3.3 (0.7) 5 (0.6) 2.2 (0.4) 2.3 (0.4) 7.2 (0.3) 4.7 (1.8) 3.3 (0.9) 6.8 (1.9) 3.8 (0.6)
Forest-VP 6.4 (0.5) 4.8 (0.6) 7 (0.4) 6.1 (0.5) 3.8 (0.5) 6.5 (0.9) 6.6 (0.8) 4.5 (0.8) 6.5 (0.8)
UTrees-kMeans 6 (0.6) 5.8 (0.5) 6.3 (0.6) 6.1 (0.6) 4.1 (0.3) 4 (0.7) 2.9 (0.6) 3.8 (1) 6 (0.7)
UTrees-KDI 5.1 (0.5) 5.1 (0.5) 5.4 (0.6) 5.6 (0.5) 4.8 (0.3) 4.5 (0.9) 4 (0.5) 3.5 (1.2) 6.4 (0.7)
UTrees 3.8 (0.5) 3.2 (0.5) 3.8 (0.4) 3.8 (0.5) 5 (0.3) 2.7 (0.6) 2.9 (0.8) 3.5 (0.8) 5.8 (0.7)

### v4

@@ 3.3 Evaluating BaltoBot on synthetic probabilistic prediction case studies
We compare our approach with Treeffuser (Beltran-Velez et al., 2024) on the “wave” synthetic dataset from Treeffuser (Beltran-Velez et al., 2024), which as shown in Figure 3 is nonlinear, multimodal, and heteroskedastic. On the raw probabilistic predictions in Figure 3(A), we see that BaltoBot and BaltoBoTabPFN are (by visual inspection) able to model the conditional distribution as well as Treeffuser. Yet this case study illustrates the two advantages of BaltoBot. First, in Figure 3(B) we show the runtime of the different methods: training, sampling, and total. To train on 5000 samples, Treeffuser took 1.1s and BaltoBot took 2.6s. But to generate 5000 samples, Treeffuser took 5.0s while BaltoBot took 0.72s, for 7$\sim$ 7$\times$ 7 × speedup. Second, BaltoBot offers the ability to estimate a closed-form probability density function (pdf) of the predictive distribution as shown in Figure 3(C); in contrast, Treeffuser can only sample from the predictive distribution.
@@
3.4 Sales forecasting with uncertainty
We employ the M5 sales forecasting Kaggle dataset (Makridakis and Howard, 2020) to compare BaltoBot with other probabilistic prediction methods. The dataset has five years of sales data from ten Walmart stores, and the task requires predicting the (heavy-tailed) number of units sold given a product’s attributes and previous sales. We use the exact same data preparation used for Treeffuser (Beltran-Velez et al., 2024) experiments, which yields 1k products, 120k training samples, and 10k test samples. As in the Treeffuser evaluation (Beltran-Velez et al., 2024), we evaluate probabilistic predictions with the continuous ranked probability score (CRPS), and evaluate the conditional mean predictions with the mean absolute error (MAE) and root mean-squared error (RMSE).
[…]
We report results in Table 5. In addition to ours’ and Treeffuser, we report results for Deep Ensembles (Lakshminarayanan et al., 2017), IBUG (Brophy and Lowd, 2022), NGBoost Poisson (Duan et al., 2020), and Quantile Regression Forests (Meinshausen and Ridgeway, 2006). For methods other than our own, we report the metrics provided in Table 2 of (Beltran-Velez et al., 2024). Overall, our proposed methods outperform previous methods at combining excellent performance on both conditional distribution prediction and conditional mean prediction. Treeffuser and BaltoBot (both with tuning) tie for first-place according to CRPS, yet BaltoBot outperforms Treeffuser on RMSE and MAE. The winners on conditional mean metrics (RMSE and MAE) are Deep Ensembles and BaltoBoTabPFN, yet BaltoBoTabPFN (no tuning) strongly outperforms Deep Ensembles on CRPS.
Table 5: Sales forecasting evaluation on M5 dataset. We highlight the best 2 methods for each metric. The best of Treeffuser versus ours (with tuning) is boldface blue; the best of Treeffuser versus ours (without tuning) is boldface brown.
Method CRPS ×$\times 10^{-1}(\downarrow)$ RMSE ×$\times 10^{0}(\downarrow)$ MAE ×$\times 10^{0}(\downarrow)$
Deep Ensembles 7.05 2.03 0.97
IBUG 8.90 2.12 1.00
NGBoost Poisson 6.86 2.33 0.99
Quantile Regression Forests 7.11 2.88 1.01
Treeffuser 6.44 2.09 0.99
BaltoBot 6.44 2.07 0.98
Treeffuser (no tuning) 6.62 2.09 0.99
BaltoBot (no tuning) 6.69 2.19 0.98
BaltoBoTabPFN (no tuning) 6.66 2.06 0.97
4 Limitations
While UnmaskingTrees is overall state-of-the-art on the tabular imputation benchmark, MissForest still outperformed on the metrics based on Wasserstein distance to train and test dataset distributions. And Forest-Flow still won on vanilla generation benchmark (without any missingness). It remains to be seen whether a single method can be developed which wins on all scenarios and metrics. While BaltoBoTabPFN performed well on probabilistic prediction tasks, when used as a subroutine in UnmaskingTabPFN, it is very slow and experienced out-of-memory errors on the (Jolicoeur-Martineau et al., 2024b) benchmark on our machine. Further improvements either to it, or to how it is employed, are needed to make it practical for all but the smallest datasets.
5 Discussion and Related Work
@@
Diffusion modeling has recently gained popularity in tabular ML (Zheng and Charoenphakdee, 2022; Jolicoeur-Martineau et al., 2024b; Beltran-Velez et al., 2024; Kotelnikov et al., 2023). Our proposed approach is an instance of the autoregressive discrete diffusion framework (Hoogeboom et al., 2021), instances of which have shown success in a variety of tasks (Yang, 2019; Austin et al., 2021; Kitouni et al., 2024; Jolicoeur-Martineau et al., 2024a). Yet our results call into question whether diffusion is beneficial for tabular conditional generation, or whether autoregression is sufficient for our setting. It has been observed that diffusion is autoregression in frequency space, progressing from low frequencies to high frequencies, which makes it a good match for image data with its power law spectra (Rissanen et al., 2022; Dieleman, 2024; Stewart, 2024). In tabular datasets without this phenomena, we would expect diffusion modeling to be less advantageous.
@@
Why is ForestDiffusion better at vanilla generative modeling, while UnmaskingTrees is better on missing data problems? We offer two speculative explanations. First, imputation is a conditional modeling scenario, except that you do not know the partition of the features into input features and output features a priori. One could address imputation by learning all possible $2^{D}$ conditional distributions, but this is impractical for large $D$, so one would prefer to learn a single joint distribution. Both autoregression and diffusion are ways of learning a joint distribution; because autoregression does so by learning conditional distributions, it is more suited to the conditional modeling imputation setting. Second, for missing data, diffusion has a train-inference gap: during training, observed features begin the reverse process from $\mathcal{N}(0,1)$; during inference for imputation, observed features begin the reverse process at their actual values. On the other hand, the advantages of diffusion modeling give it superiority when these problems can be avoided.
@@
We also note that MissForest (Stekhoven and Bühlmann, 2012), hailing from statistical literature on multiple imputation, has yet to be completely dethroned. Future progress in tabular conditional generation may require going back to the well of this traditional literature. As one example, we observe that MissForest exploits feature missingness fraction information, but we are not aware of any “machine learning” approaches which do so. The statistical literature has also previously explored the value of conditional modeling for joint modeling (Gelman and Raghunathan, 2001; Liu et al., 2014; Kropko et al., 2014). Indeed, our UnmaskingTrees approach, and all autoregressive modeling, is presaged by the full-mechanism bootstrap (Efron, 1994).
Finally, we observe where randomness enters into our generation process compared to previous work. Flow-matching (Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2022) (used in Forest-Flow) injects randomness solely at the beginning of the reverse process via Gaussian sampling, whereas diffusion modeling (Sohl-Dickstein et al., 2015; Song and Ermon, 2019) (used in Forest-VP) injects randomness both at the beginning and during the reverse process. In contrast, because our method starts with a fully-masked sample, it injects randomness gradually during the generation process. First, we randomly generate the order over features for unmasking. Second, we do not “greedily decode” to the most likely leaf in the meta-tree, but instead sample according to predicted probabilities. Third, for continuous features, having sampled a particular meta-tree leaf bin, we sample from within the bin, treating it as a uniform distribution.
@@ 6 Conclusions
We proposed tree-based autoregressive modeling of tabular data, especially for data with missingness. For the subproblem of conditional probabilistic prediction of individual variables, we presented a hierarchical partitioning method with benefits over vanilla quantization and diffusion-based probabilistic prediction. We then considered each of these as meta-algorithms that enable pure in-context learning-based modeling using TabPFN as base classifier. We showed SotA results for imputation and for generation given data with missingness, and on probabilistic prediction for sales forecasting.
Appendix A Algorithmic descriptions of UnmaskingTrees and BaltoBot
The training algorithm for UnmaskingTrees is given in 1. The training and inference algorithms for BaltoBot are given in 2 and 3, respectively.
Algorithm 1 Unmasking Trees training
0:  dataset ${\mathbf{X}}\in{\mathbb{R}}^{N\times D}$; duplication factor $K$.
1:  {# Build self-supervised training set}
2:  Set ${\mathbf{X}}_{\textrm{train}}=\emptyset$, ${\mathbf{Y}}_{\textrm{train}}=\emptyset$.
3:  for $k=1,\dots,K$ do
4:     for $n=1,\dots,N$ do
5:        Draw random permutation $\sigma$ from ${\mathcal{U}}(G_{D})$
6:        Set $\bm{x}:={\mathbf{X}}_{n,:}$ and $\bm{y}:={\mathbf{X}}_{n,:}$.
7:        for $d=1,\dots,D$ do
[…]
9:           Append ${\mathbf{X}}_{\textrm{train}}:={\mathbf{X}}_{\textrm{train}}\cup\{\bm{x}\}$, ${\mathbf{Y}}_{\textrm{train}}:={\mathbf{Y}}_{\textrm{train}}\cup\{\bm{y}\}$
10:        end for
11:     end for
12:  end for
13:  {# Train conditional generation models}
14:  for $d=1,\dots,D$ do
15:     if feature $d$ in ${\mathbf{X}}$ is a continuous feature then
16:        Run BaltoBot with $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
17:     else
18:        Train XGBClassifier on $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
19:     end if
20:  end for
Algorithm 2 BaltoBot training
0:  dataset $({\mathbf{X}}\in{\mathbb{R}}^{N\times D},{\mathbf{y}}\in{\mathbb{R}}^{N})$; BaltoBot meta-tree height $H$;
1:  if H = 0 or unique$({\mathbf{y}})=C$ for some constant $C$ then
2:     Save $\texttt{bounds}:=(\min({\mathbf{y}}),\max({\mathbf{y}}))$.
3:  else
4:     Obtain split point $p$ from KDI quantization on ${\mathbf{y}}$.
5:     Train XGBoost binary classifier on $({\mathbf{X}},\bm{1}\{{\mathbf{y}}\leq p\})$.
6:     Train “left-child” BaltoBot on $\{({\mathbf{X}}^{(i)},{\mathbf{y}}^{(i)})\in({\mathbf{X}},{\mathbf{y}})|{\mathbf{y}}^{(i)}\leq p\}$, with height $H-1$.
7:     Train “right-child” BaltoBot on $\{({\mathbf{X}}^{(i)},{\mathbf{y}}^{(i)})\in({\mathbf{X}},{\mathbf{y}})|{\mathbf{y}}^{(i)}>p\}$, with height $H-1$.
8:  end if
Algorithm 3 BaltoBot inference
0:  input query ${\mathbf{x}}\in{\mathbb{R}}^{D}$; trained BaltoBot model.
1:  if bounds is defined then
2:     Sample uniformly from $U(\texttt{bounds})$.
3:     Return.
4:  else
5:     Obtain prediction from XGBoost binary classifier.
6:     if prediction = left-child then
7:        Run inference on “left-child” BaltoBot with input query ${\mathbf{x}}$.
8:     else if prediction = right-child then
9:        Run inference on “right-child” BaltoBot with input query ${\mathbf{x}}$.
10:     end if
11:  end if
Appendix B Ablation experiment with imputation - raw scores
Raw scores (shown in Table 6) demonstrate that UnmaskingTrees on its own improves upon Forest-VP’s diffusion approach. We also see that KDI quantization (with 20 bins) contributes to improvement beyond k-Means (also 20 bins), and that BaltoBot yields even further improvement.